A large department store found that it averages 362 customers per hour. Assume that the standard...

  1. A large department store found that it averages 362 customers per hour. Assume that the standard deviation is 29.6 and a random sample of 40 hours was used to determine the average. Find the 98% confidence interval of the population mean.
Answer #1

Solution :

Given that,

= 362

= 29.6

n = 40

At 98% confidence level the z is ,

  = 1 - 98% = 1 - 0.98 = 0.02

/ 2 = 0.02 / 2 = 0.01

Z/2 = Z0.01 = 2.326

Margin of error = E = Z/2* ( /n)

= 2.326 * (29.6 / 40)

= 10.8861

At 98% confidence interval estimate of the population mean is,

- E < < + E

362 - 10.8861 < < 362 + 10.8861

351.1139< < 372.8861

(351.1139 ,   372.8861)